OpenStudy (cicilybailey14):

Help! Do the ordered pairs below represent a relation, a function, both a relation and a function, or neither a relation nor a function? (-3,0) , (0,-3) , (7,-10) , (9,-12) A. function only B. relation only C. both a relation and a function D. neither a relation nor a function

OpenStudy (anonymous):

point 1 (-3,0) when y=0, x=-3 point 2 (0,-3) y=-3 when x=0 y=mx+b from second point b=-3 y=mx-3 from 1 point 0=m(-3)-3 ......m=-1 y=-x-3

OpenStudy (anonymous):

now use this formula and see if it get the numbers.... 1 ( -3,0) it works 2 (0,-3) it works 3 (7,-10).....y=-(7) -3=-10 it works 4 (9,-12).......y=-(9)-3= -12 it works it is a function

OpenStudy (solomonzelman):

Don't even know which way to do it. Will choose the easiest way: Suppose: \(\large\color{slate}{ a_{-3}=0 }\) \(\large\color{slate}{ a_0=-3 }\) \(\large\color{slate}{ a_7=-10 }\) \(\large\color{slate}{ a_9=-12 }\) now, starting from an index n=1 \(\large\color{slate}{ a_{1}=0 }\) \(\large\color{slate}{ a_4=-3 }\) \(\large\color{slate}{ a_{11}=-10 }\) \(\large\color{slate}{ a_{13}=-12 }\) the differences b/w: \(\large\color{slate}{ a_{1},~a_4 }\) \(\large\color{teal}{ a_{4}=a_1+d(n-1) }\) \(\large\color{teal}{ -3=0+d(4-1) }\) \(\large\color{teal}{ -3=3d }\) \(\large\color{teal}{ d=1 }\) \(\large\color{slate}{ a_{4},~a_{11} }\) \(\large\color{teal}{ a_{11}=a_{4}+d(7-1) }\) \(\large\color{teal}{ -10=-3+d(8-1) }\) \(\large\color{teal}{ -7=7d }\) \(\large\color{teal}{ d=-1 }\) \(\large\color{slate}{ a_{11},~a_{13} }\) \(\large\color{teal}{ a_{13}=a_{11}+d(3-1) }\) \(\large\color{teal}{ -12=-10+d(3-1) }\) \(\large\color{teal}{ -2=2d }\) \(\large\color{teal}{ d=-1 }\)